# The Fall of ROME: Understanding the Collapse of LLMs in Model Editing

### v1

@@ Abstract
Despite significant progress in model editing methods, their application in real-world scenarios remains challenging as they often cause large language models (LLMs) to collapse. Among them, ROME is particularly concerning, as it could disrupt LLMs with only a single edit. In this paper, we study the root causes of such collapse. Through extensive analysis, we identify two primary factors that contribute to the collapse: i) inconsistent handling of prefixed and unprefixed keys in the parameter update equation may result in very small denominators, causing excessively large parameter updates; ii) the subject of collapse cases is usually the first token, whose unprefixed key distribution significantly differs from the prefixed key distribution in autoregressive transformers, causing the aforementioned issue to materialize. To validate our analysis, we propose a simple yet effective approach: uniformly using prefixed keys during editing phase and adding prefixes during the testing phase. The experimental results show that the proposed solution can prevent model collapse while maintaining the effectiveness of the edits [^1] The code will be released after the review process ends..
The Fall of ROME: Understanding the Collapse of LLMs in Model Editing
Wanli Yang  Fei Sun [^2] Corresponding author.  Jiajun Tan   Xinyu Ma   Du Su   Dawei Yin   Huawei Shen♠♡ CAS Key Laboratory of AI Safety, Institute of Computing Technology, CAS University of Chinese Academy of Sciences       Baidu Inc. yangyywl@gmail.com    sunfei@ict.ac.cn
@@ 1 Introduction
Intuitively, for a knowledge tuple (subject, relation, object), ROME takes a prompt constructed from the subject and relation as input and models the knowlege in a key-value format. Here, the key is a vector representation of the subject, and the value is a vector representation that can produce the target object, obtained by transforming the key through a transformation matrix. To insert a new fact about a subject, ROME adjusts the transformation matrix to matche the subject’s key vector with the new fact’s value vector, as described in Eq. 3.
To uncover the underlying causes, we investigate the differences in parameter update process of ROME between collapse cases (i.e., samples that induce collapse) and normal cases (i.e., samples that do not). The results reveal that the collapse directly stems from the anomalously small denominator within the parameter update equation in Eq. 3. This anomaly originates from the irregular implementation of the keys in the denominator, where one key is derived with varying prefixes (prefixed key) and another without any prefix (unprefixed key). This issue has also been independently identified by Gupta et al. (2024a) simultaneously. However, it is still unclear why the irregular implementation only leads to collapse in collapse cases.
To answer this question, we examine the distribution of elements in the denominator. We observe that, in collapse cases, the distribution of the unprefixed keys exhibits significant differences from the prefixed keys. This leads to an exceptionally small denominator in the update equation, which in turn causes the model to collapse.
To elucidate the anomalous behavior observed in collapse cases, we conduct an analysis starting from their characteristics. The collapse cases of GPT-2-XL Radford et al. (2019) and GPT-J Wang and Komatsuzaki (2021) exhibit a consistent pattern: the subjects in nearly all of these instances correspond to the first tokens within their respective prompts. Furthermore, we discover that the representation distribution of the first tokens markedly diverges from that of the subsequent tokens in these autoregressive models. These two factors, working in concert, lead to the anomalous distribution of unprefixed keys in collapse cases.
To validate our findings, we propose unifying all keys as prefixed during editing to prevent model collapse. When using the edited model, we prepend a random text prefix for instances where subjects are in the first token to ensure consistency with the editing process. Experiments validate that our proposed method effectively prevents model collapse while ensuring the success of edits.
@@
  • We perform comprehensive analyses to identify two factors behind ROME’s collapse: i) inconsistent implementation of key vectors; ii) anomalous distribution of first token representations.
  • We propose a straightforward solution to prevent collapse while maintaining editing efficacy.
@@ 2 Background
ROME Meng et al. (2022) hypothesizes that MLP modules in Transformer architecture Vaswani et al. (2017) can be modeled as a linear key-value associative memory Geva et al. (2021). Under this hypothesis, a knowledge triplet $(s,r,o)$ corresponds to a key-value pair $(k,v)$, where $k$ represents the subject $s$, and $v$ encodes the property $(r,o)$ for $s$. The entire knowledge within a model can thus be represented as a set of key vectors $K=[k_{1},\cdots,k_{n}]$ and value vectors $V=[v_{1},\cdots,v_{n}]$. A linear operation $W$ matches keys to values by solving $WK\approx V$. In practice, for two-layer MLP in a specific transformer block determined by a Causal Tracing mechanism Meng et al. (2022), outputs of the first layer form a key $k$, and the second layer (parameterized with $W$) retrieves an associated value $v$ based on this key $k$.
In this context, to replace the current knowledge $(s,r,o)$ with a new knowledge tuple $t^{*}=(s,r,o^{*})$, we need to find the corresponding key vector $k_{*}$ and the new value vector $v_{*}$. To simulate various contexts for generalization, ROME assigns $k_{*}$ as an average vector derived from subject $s$ with a small set of $N$ randomly sampled prefixes:
$k_{*}=\frac{1}{N}\sum_{i=1}^{N}\mathcal{K}\left(x_{i}\oplus s\right),$ (1)
@@
To illustrate the selection of $v_{*}$, we take the subject $s{=}$ United States and relation $r{=}$ president of as an example. A specifically designed loss function is utilized to optimize $v_{*}$ so that it can produce $o^{*}=$ Joe Biden when provided with the prompt $\mathtt{p}(s,r)=$ The president of the United States is.
Given the computed $(k_{*},v_{*})$, ROME finds optimal $\widehat{W}$ to solve the following problem:
$\underset{\widehat{W}}{\arg\min}\|\widehat{W}K-V\|\ \text{ subject to }\widehat{W}k_{*}=v_{*}$ (2)
@@
$\widehat{W}=W+\underbrace{\frac{\left(v_{*}-Wk_{*}\right)\left(C^{-1}k_{*}\right)^{\top}}{\left(C^{-1}k_{*}\right)^{\top}k_{*}}}_{\text{update matrix }\Delta}$ (3)
where $W$ denotes the weight matrix of the second layer of the MLP before editing, $\widehat{W}$ denotes the weight matrix after editing, and $C{=}KK^{\top}$ is a pre-cached constant. Interested readers are directed to Meng et al. (2022) for a detailed introduction.
@@ 3 Why Does ROME Cause Collapse?
Previous studies Yang et al. (2024); Gupta et al. (2024b) have revealed that a single edit of ROME can induce LLMs to collapse. To further analyze the cause, we investigate the differences in parameter updates between samples that induce collapse and those that do not. For this purpose, we introduce two distinct subsets: i) collapse cases, using the HardCF set built by Yang et al. (2024), which includes collapse cases on GPT-2-XL, GPT-J, and Llama2-7b from the COUNTERFACT dataset Meng et al. (2022); and ii) normal cases, comprising 1000 random samples from the remaining part of COUNTERFACT.
@@ 3.1 Inconsistent Keys in Editing
Component Cases GPT-2-XL GPT-J Llama2-7b numerator: collapse $168.55$ $140.27$ $4.57$ $\left(v_{*}-Wk_{*}\right)\left(C^{-1}k_{*}\right)^{\top}$ normal $79.91$ $88.69$ $16.52$ denominator: collapse $0.04$ $0.04$ $0.01$ $\left(C^{-1}k_{*}\right)^{\top}k_{*}$ normal $9.60$ $12.78$ $2.63$
@@
These results guide our focus to the denominator $(C^{-1}k_{*})^{\top}k_{*}$. Given that the matrix $C$ is a constant for both collapse cases and normal cases, our analysis is primarily focused on the key $k_{*}$. We revisited the official implementation of ROME and identified that different variants of $k_{*}$ are used. Specifically, only $k_{*}$ within $(C^{-1}k_{*})^{\top}$ is the prefixed key as in Eq. 1. In contrast, $k_{*}$ in other positions is unprefixed, utilizing a representation over the subject $s$ without any prefix, denoted as $k_{*}^{s}=\mathcal{K}\left(s\right)$. However, ideally, all key $k_{*}$ in Eq. 3 should be the same vector, i.e., the average representation derived from a set of prefixed subjects as in Eq. 1.
To verify if this inconsistency of keys is responsible for the collapse, we substitute all $k_{*}^{s}$ with $k_{*}$ in the implementation. The aligned implementation is referred to as Consistent-ROME, C-ROME for short. We evaluate the different implementations on collapse and normal cases using perplexity on the ME-PPL50 dataset, whose effectiveness has been validated by Yang et al. (2024). According to Table 2, C-ROME with aligned implementation of $k_{*}$ does not significantly alter the edited models, avoiding the sharp increase in perplexity seen with ROME. This demonstrates that such inconsistency of $k_{*}$ in the update matrix $\Delta$ is a primary factor behind ROME-induced model collapse.
@@ 3.2 Anomalous Key Distribution for Collapse
While unifying the keys as $k_{*}$ can prevent model collapse, it remains unclear why inconsistent keys only encounter issues in collapse cases.
To enhance intuitive understanding, we analyze the spatial distribution of $C^{-1}k_{*}$ and $k_{*}^{s}$ in the denominator for different cases by projecting them into a two-dimensional space using t-SNE Van der Maaten and Hinton (2008). Taking the results of GPT-2-XL in Figure 1(a) as an example, in normal cases, the distribution of $C^{-1}k_{*}$ and $k_{*}^{s}$ show no significant differences. However, a noticeable divergence in the distribution occurs in collapse cases, explaining the exceptionally small denominators.
Considering that $C$ is a constant, the distinctions between normal and collapse cases should arise from the variations in the prefixed key $k_{*}$ and the unprefixed key $k_{*}^{s}$. Figure 1(b) clearly illustrates that the distribution of $k_{*}^{s}$ in collapse cases significantly diverge from those of $k_{*}$. This confirms that in collapse cases, the significant differences between $k_{*}$ and $k_{*}^{s}$ result in a particularly small denominator in the update matrix, which in turn leads to the collapse of the edited model. Similar phenomena are also observed in other LLMs, detailed in § A.1.
@@ 3.3 Special Role of the First Token
To elucidate the anomalous distribution of $k_{*}^{s}$ in collapse cases, we focus our analysis on their characteristics. We observed a common pattern in the collapse cases for both GPT-2-XL and GPT-J: in almost all instances, the subjects consist of a single word, which is encoded as a single token and positioned at the beginning of the input prompt $\mathtt{p}(s,r)$ [^2] The only exception involves few instances with subjects like “Jackson Jackson” in the collapse cases of GPT-J.. Therefore, the unprefixed key $k_{*}^{s}$ for a collapse case is the intermediate representation within the MLP layer of the first token in the input prompt. This inspires us to investigate whether the anomalous distribution of $k_{*}^{s}$ in collapse cases can be attributed to their place as the first tokens in the prompts.
@@
To explore this, we first examined the representation distribution of the first tokens in the prompts for normal cases. The results presented in Figure 2(a) indicate that, within GPT-2-XL, the first tokens of normal cases consistently exhibit an abnormal distribution similar to that of $k_{*}^{s}$ in collapse cases. From an opposing perspective, to verify whether artificially shifting the $k_{*}^{s}$ in collapse cases away from the first position would eliminate the anomaly in distribution, we prefixed the prompts of collapse cases with randomly sampled texts. This adjustment results in their distribution aligning with that of normal cases, as illustrated in Figure 2(b). The results suggest that the anomalous distribution of $k_{*}^{s}$ for collapse cases in ROME is not related to the editing process. Instead, it is due to the unique pattern of their subjects encountering the special distribution of the first token in GPT-like models.
We speculate that this phenomenon arises from the inherent nature of autoregressive models, where the first token cannot interact with any other token except itself. As a counterexample with non-autoregressive architecture, the representation distribution of first tokens in T5-3B encoder Raffel et al. (2020) does not differ from that of subsequent tokens. This may be attributed to the bidirectional attention in the encoder, which enables interactions between the first token and subsequent tokens. A detailed analysis is presented in Appendix A.2.
@@
It is important to note that Llama2-7b Touvron et al. (2023) avoids collapse in such cases due to its tokenizer incorporating a special token, <s>, at the beginning of the encoding process, which shifts the subject from being the first token. In fact, we found that Llama2-7b also succumbs to collapse when the special token <s> is not prepended, with results detailed in Appendix A.3.
@@
Figure 2: t-SNE visualization of representation distributions of (a) the first token in randomly sampled normal prompts; (b) $k_{*}^{s}$ in prefixed collapse prompts.

### v2

# Understanding the Collapse of LLMs in Model Editing

@@ Abstract
Despite significant progress in model editing methods, their application in real-world scenarios remains challenging as they often cause large language models (LLMs) to collapse. Among them, ROME is particularly concerning, as it could disrupt LLMs with only a single edit. In this paper, we study the root causes of such collapse. Through extensive analysis, we identify two primary factors that contribute to the collapse: i) inconsistent handling of prefixed and unprefixed keysin the parameter update equation may result in very small denominators, causing excessively large parameter updates; ii) the subject of collapse cases is usually the first token, whose unprefixed key distribution significantly differs from the prefixed key distribution in autoregressive transformers, causing the aforementioned issue to materialize. To validate our findings, we propose a simple yet effective approach: uniformly using prefixed keys during editing phase and adding prefixes during testing phase to ensure the consistency between training and testing. The experimental results show that the proposed solution can prevent model collapse while maintaining the effectiveness of the edits [^1] Code and data are available at: https://github.com/WanliYoung/Collapse-in-Model-Editing..
Understanding the Collapse of LLMs in Model Editing
Wanli Yang♠♡    Fei Sun [^2] Corresponding author. Jiajun TanXinyu MaDu SuDawei YinHuawei Shen♠♡ CAS Key Laboratory of AI Safety, Institute of Computing Technology, CAS University of Chinese Academy of Sciences       Baidu Inc. yangwanli24z@ict.ac.cn     sunfei@ict.ac.cn
@@ 1 Introduction
Intuitively, for a knowledge tuple (subject, relation, object), ROME takes a prompt constructed from the subject and relation as input and models the knowlege in a key-value format. Here, the key is a vector representation of the subject within the prompt, and the value is a vector representation capable of yielding the target object, obtained by transforming the key through a transformation matrix. To insert a new fact about a subject, ROME adjusts the transformation matrix to match the key of the subject with the value of the new fact, as described in Eq. 3.
To uncover the underlying causes of ROME’s collapse, we investigate the differences in parameter update process of ROME between collapse cases (i.e., samples that induce collapse) and normal cases (i.e., samples that do not). The results reveal that the collapse directly stems from the anomalously small denominator within the parameter update equation (Eq. 3). This anomaly originates from the irregular implementation of the keys in the denominator, where one is derived by prepending varying prefixes to the subject to simulate diverse contexts (termed prefixed key), while the other is obtained directly from the original subject without any prefix (termed unprefixed key). This issue has also been independently identified by Gupta et al. (2024a) concurrently. However, it is still unclear why the irregular implementation only fails in collapse cases.
To answer this question, we examine the distribution of elements in the denominator. It reveals that, in collapse cases, the distribution of the unprefixed keys exhibits significant difference from the prefixed keys. This leads to an exceptionally small denominator in the update equation, which in turn causes the model to collapse.
Figure 1: To update “the president of the United States” from “Donald Trump” to “Joe Biden”, ROME locates the knowledge into the MLP module within a specific transformer block using the Causal Tracing mechanism. It then adjusts the second layer of MLP (i.e., weight matrix $W$) to change the value $\bm{v}$ for the key $\bm{k}$ that represents the subject “the United States” to a new value $\bm{v}_{*}$, thereby inducing the LLMs to predict the target object “Joe Biden”.
To elucidate the anomalous behavior observed in the collapse cases, we conduct an analysis starting from their characteristics. The collapse cases of both GPT-2-XL Radford et al. (2019) and GPT-J Wang and Komatsuzaki (2021) exhibit a consistent pattern: the subjects in nearly all of these instances correspond to the first tokens within their respective prompts. Furthermore, we discover that the representation distribution of the first tokens markedly diverges from that of the subsequent tokens in these autoregressive models. These two factors, working in concert, lead to the anomalous distribution of unprefixed keys in collapse cases.
To validate our findings, we propose unifying all keys as prefixed during editing to prevent model collapse. To ensure consistency with the editing process, when using the edited model, we prefix a random text for instances where subjects are in the first token. Experiments validate that our proposed method effectively prevents model collapse while ensuring the success of edits.
@@
  • Comprehensive analysis that identifies two factors behind ROME’s collapse: i) inconsistent implementation of key vectors; ii) anomalous distribution of first token representations.
  • A straightforward solution to prevent collapse while maintaining editing efficacy.
@@ 2 Background
ROME Meng et al. (2022) hypothesizes that the MLP modules in the Transformer architecture Vaswani et al. (2017) can be modeled as a linear key-value associative memory. Under the hypothesis in ROME, a knowledge triplet $(s,r,o)$ corresponds to a key-value pair $(\bm{k},\bm{v})$, where $\bm{k}$ represents the subject $s$, and $\bm{v}$ encodes the property $(r,o)$ for $s$. The entire knowledge within a model can thus be represented as a set of key vectors $K=[\bm{k}_{1},\dots,\bm{k}_{n}]$ and value vectors $V=[\bm{v}_{1},\dots,\bm{v}_{n}]$. A linear operation $W$ matches keys to values by solving $WK\approx V$.
In practice, for an input prompt $\mathtt{p}(s,r)$, the recall of the target object $o$ mainly occurs within a two-layer MLP in a specific transformer block identified by the Causal Tracing mechanism Meng et al. (2022). Specifically, output of the first layer for the subject $s$ forms a key $\bm{k}$, and the second layer (parameterized with $W$) retrieves an associated value $\bm{v}$ based on this key $\bm{k}$, ultimately inducing the LLMs to predict the target object $o$.
In this context, to replace the current knowledge $(s,r,o)$ with a new knowledge tuple $t^{*}=(s,r,o^{*})$, we need to find the corresponding key $\bm{k}$ and the new value $\bm{v}_{*}$. To simulate various contexts for generalization, ROME assigns $\bm{k}$ as an average vector $\overline{\bm{k}}$ derived from subject $s$ with a small set of $N$ randomly sampled prefixes:
$\overline{\bm{k}}=\frac{1}{N}\sum_{i=1}^{N}\mathcal{K}\left(x_{i}\oplus s\right)$ (1)
@@
To illustrate the selection of $\bm{v}_{*}$, we take the subject $s{=}$United States” and relation $r{=}$president of” as an example. A specifically designed loss function is utilized to optimize $\bm{v}_{*}$ so that it can produce $o^{*}=$Joe Biden” when given the prompt $\mathtt{p}(s,r)=$The president of the United States is”.
With the computed $(\overline{\bm{k}},\bm{v}_{*})$, ROME finds optimal $\widehat{W}$ by solving the following problem:
$\underset{\widehat{W}}{\arg\min}\|\widehat{W}K-V\|\ \text{ subject to }\widehat{W}\overline{\bm{k}}=\bm{v}_{*}$ (2)
@@
$\widehat{W}=W+\underbrace{\frac{\left(\bm{v}_{*}-W\overline{\bm{k}}\right)\left(C^{-1}\overline{\bm{k}}\right)^{\top}}{\left(C^{-1}\overline{\bm{k}}\right)^{\top}\overline{\bm{k}}}}_{\text{update matrix }\Delta}$ (3)
where $W$ denotes the weight matrix of the second layer in the MLP before editing, $\widehat{W}$ denotes the weight matrix after editing, and $C{=}KK^{\top}$ is a pre-cached constant.
The complete editing process of ROME is illustrated in Figure 1. Interested readers are directed to Meng et al. (2022) for a detailed introduction.
@@ 3 Why Does ROME Cause Collapse?
Previous studies Yang et al. (2024); Gupta et al. (2024b) have revealed that a single edit of ROME can induce LLMs to collapse. To further analyze the cause, we investigate the differences in parameter updates between samples that induce collapse and those do not. For this purpose, we introduce two distinct subsets: i) collapse cases, using the HardCF set built by Yang et al. (2024), which includes collapse cases on GPT-2-XL, GPT-J, and Llama2-7b from the COUNTERFACT dataset Meng et al. (2022); and ii) normal cases, comprising 1000 random samples from the remaining part of COUNTERFACT.
@@ 3.1 Inconsistent Keys in Editing
Component Cases GPT-2-XL GPT-J Llama2-7b numerator: collapse $168.55$ $140.27$ $4.57$ $\left(\bm{v}_{*}-W\overline{\bm{k}}\right)\left(C^{-1}\overline{\bm{k}}\right)^{\top}$ normal $79.91$ $88.69$ $16.52$ denominator: collapse $0.04$ $0.04$ $0.01$ $\left(C^{-1}\overline{\bm{k}}\right)^{\top}\overline{\bm{k}}$ normal $9.60$ $12.78$ $2.63$
@@
These results guide our focus to the key $\overline{\bm{k}}$ in the denominator $(C^{-1}\overline{\bm{k}})^{\top}\overline{\bm{k}}$, given that the matrix $C$ is a constant for both collapse cases and normal cases. We revisited the official implementation of ROME and identified that different variants of $\overline{\bm{k}}$ are used. Specifically, only $\overline{\bm{k}}$ within $(C^{-1}\overline{\bm{k}})^{\top}$ is the prefixed key as in Eq. 1. In contrast, $\overline{\bm{k}}$ in other positions is unprefixed, utilizing a representation over the subject $s$ without any prefix, denoted as $\bm{k}^{u}=\mathcal{K}\left(s\right)$. However, ideally, all $\overline{\bm{k}}$ in Eq. 3 should be the same, i.e., the average representation derived from a set of prefixed subjects as in Eq. 1.
To verify if this inconsistency of keys is responsible for the collapse, we substitute all $\bm{k}^{u}$ with $\overline{\bm{k}}$ in the implementation. The aligned implementation is referred to as Consistent-ROME, C-ROME for short. We evaluate the different implementations on collapse and normal cases using perplexity on the ME-PPL50 dataset, whose effectiveness has been validated by Yang et al. (2024). According to Table 2, C-ROME with aligned implementation of $\overline{\bm{k}}$ does not significantly alter the edited models, avoiding the sharp increase in perplexity seen with ROME. This demonstrates that such inconsistency of $\overline{\bm{k}}$ in the update matrix $\Delta$ is a primary factor behind ROME-induced model collapse.
@@ 3.2 Anomalous Key Distribution for Collapse
While unifying the keys as $\overline{\bm{k}}$ can prevent model collapse, it remains unclear why inconsistent keys only encounter issues in collapse cases.
To enhance intuitive understanding, we analyze the spatial distribution of $C^{-1}\overline{\bm{k}}$ and $\bm{k}^{u}$ in the denominator for different cases by projecting them into a two-dimensional space using t-SNE Van der Maaten and Hinton (2008). Taking the results of GPT-2-XL in Figure 2(a) as an example, in normal cases, the distribution of $C^{-1}\overline{\bm{k}}$ and $\bm{k}^{u}$ show no significant differences. However, a noticeable divergence in the distribution occurs in collapse cases, explaining the exceptionally small denominators.
Considering that $C$ is a constant, the differences between normal and collapse cases should arise from the variations in the prefixed key $\overline{\bm{k}}$ and the unprefixed key $\bm{k}^{u}$. Figure 2(b) clearly illustrates that the distribution of $\bm{k}^{u}$ in collapse cases significantly diverge from those of $\overline{\bm{k}}$. This confirms that in collapse cases, the significant differences between $\overline{\bm{k}}$ and $\bm{k}^{u}$ result in a particularly small denominator in the update matrix, which in turn leads to the collapse of the edited model. Similar phenomena are also observed in other LLMs, detailed in § A.1.
@@ 3.3 Special Role of the First Token
To elucidate the anomalous distribution of $\bm{k}^{u}$ in collapse cases, we focus our analysis on their characteristics. A common pattern is observed in the collapse cases for both GPT-2-XL and GPT-J: in almost all instances, the subjects consist of a single word, which is encoded as a single token and positioned at the beginning of the input prompt $\mathtt{p}(s,r)$ [^2] The only exception involves few instances with subjects like “Jackson Jackson” in the collapse cases of GPT-J.. Therefore, the unprefixed key $\bm{k}^{u}$ for a collapse case is the intermediate representation within the MLP layer of the first token in the input. This inspires us to investigate whether the anomalous distribution of $\bm{k}^{u}$ in collapse cases can be attributed to their position as the first tokens in the prompts.
@@
To explore this, we first examined the representation distribution of the first tokens in the prompts for normal cases. The results presented in Figure 3(a) indicate that, within GPT-2-XL, the first tokens of normal cases consistently exhibit an abnormal distribution similar to that of $\bm{k}^{u}$ in collapse cases. From an opposing perspective, to verify whether artificially shifting the $\bm{k}^{u}$ in collapse cases away from the first position would eliminate the anomaly in distribution, we prefixed the prompts of collapse cases with randomly sampled texts. This adjustment results in their distribution aligning with that of normal cases, as illustrated in Figure 3(b). These findings suggest that the anomalous distribution of $\bm{k}^{u}$ for collapse cases in ROME is not related to the editing process. Instead, it is due to the unique pattern of their subjects encountering the special distribution of the first token in GPT-2-XL and GPT-J models.
@@
It is important to note that Llama2-7b Touvron et al. (2023), Mistral-7b Jiang et al. (2023), and Llama3-8b Meta (2024) avoid collapse in such cases due to their tokenizers additionally incorporating a special token, e.g., <s>, at the beginning of the input, which shifts the subject from being the first token. In fact, we found they also succumb to collapse when the special token is removed, with results detailed in Appendix A.2.
@@
Figure 3: t-SNE visualization of representation distributions of (a) the first token in randomly sampled normal prompts; (b) $\bm{k}^{u}$ in prefixed collapse prompts.
Analysis. To elucidate the underlying reasons for the anomalous distribution of the first token in autoregressive language models, we explored two potential factors as follows.
Firstly, we speculate that this phenomenon may arise from the inherent nature of autoregressive models, where the first token cannot interact with any other token except itself. As a counterexample with non-autoregressive architecture, the representation distribution of the first token in T5-3B encoder Raffel et al. (2020) does not differ from that of subsequent tokens. This may be attributed to the bidirectional attention in the encoder, which enables interactions between the first token and subsequent tokens. A detailed analysis is presented in Appendix A.3.
Secondly, considering the specificity of the first token may originate from its position embedding, we verify it from two aspects. For collapse cases where the subjects are the first tokens, setting the position embedding of the first token as that of the second token can not completely eliminate collapse. While for normal cases where the subjects are the second tokens, replicating the position embedding of the first token onto the second token does not consistently lead to collapse. These findings suggest that while position embedding plays a role, it is not the only determining factor. The detailed investigation is provided in Appendix A.4.
Additionally, we observed that in GPT-2-XL and GPT-J, the representations of the first tokens rapidly become significantly more concentrated than those of subsequent tokens as the layers progress. However, this phenomenon does not appear in Llama2-7b, Mistral-7b, and Llama3-8b. A detailed investigation is presented in Appendix A.5.
Model GPT-2-XL GPT-J Mistral-7b Llama3-8b Ori PPL $68.39$ $50.34$ $51.75$ $41.67$ Max PPL $68.91$ $50.59$ $52.19$ $43.98$
Table 3: The maximum perplexity for various LLMs edited by ROME on the collapse cases of Llama2-7b, with their original perplexity for comparison.
Regarding the collapse cases of Llama2-7b, we found that the subjects of them terminate with a period “.”. It is worth noting that, such cases are extremely rare, amounting to just $21$ out of $21919$ samples in the COUNTERFACT dataset. Furthermore, they do not induce model collapse in various other models, including GPT-2-XL, GPT-J, Mistral-7b and Llama3-8b (the successor of Llama2-7b), as shown in Table 3. Consequently, we have decided not to pursue an exhaustive investigation of this isolated phenomenon.